OpenStudy (anonymous):

last one like this....find n. x^n ---- = x^3 x^7

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^n}{x^7}=x^3\] \[\Large x^{n-7}=x^3\] \[\Large n-7=3\] \[\Large n-7+7=3+7\] \[\Large n=10\]

OpenStudy (anonymous):

thanks(:

jimthompson5910 (jim_thompson5910):

yw, glad to be of help